OpenStudy (anonymous):

Simplify the given expression: 4 divided by the quantity of 3 minus 2i HELP!!!! MEDAL!!!!

OpenStudy (anonymous):

you mean \[\frac{ 4 }{ 3-2i }\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I'm not going to do the exercise for you. You can only learn when working together

OpenStudy (anonymous):

YES AND I WILL

OpenStudy (anonymous):

???

OpenStudy (anonymous):

So my question is whether what you need to simplify looks like what I wrote in my first reply

OpenStudy (anonymous):

YES it is

OpenStudy (anonymous):

Ok. What does i stand for?

OpenStudy (anonymous):

do not say just says simplify

OpenStudy (anonymous):

i is the complex number... i²=-1 and so i is the square of -1. It is not a variable. It is the symbol for something that does not exist with the real numbers. In the Real numbers you cannot take the square of a negative number. But with the Complex numbers you can because -1 = i² and you can always write the square of i² as i. This means that if you were to have to take the square of -9 you could also write it as 9.(-1) or as 9i². You can take the square from that, which will be 3i. So all these simplify questions have as a purpose to teach you how to work with the complex number i and i² and -1. Do you understand now what i means?

OpenStudy (anonymous):

I understand

OpenStudy (anonymous):

Ok so we could write the exercise in another way so that you can understand why you need to simplify it, and perhaps recognize how you need to simplify it. \[\frac{ 4 }{ 3-2i }=\frac{ 4 }{ 3-2\sqrt{-1} }\] Have you learned what you should do if there is a square root in the denominator?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

when you have a squareroot you want to get rid of it in the denominator. Instead it will appear in the numenator, but that's less bad than in the denominator. Example: Let's say you have \[\frac{ 5 }{ ? }\] You can "simplify" it (that means getting rid of the square root in the denominator) by doing this: \[\frac{ 5 }{ \sqrt{2} }\frac{ \sqrt{2} }{ \sqrt{2} }=\frac{5\sqrt{2}}{2}\] Multiplying a square root of a number with itself cancels out the square root. Example two: \[\frac{ 5 }{ 1+\sqrt{2} }\] You "simplify" it by doing this \[\frac{ 5 }{ 1+\sqrt{2} }\frac{ 1-\sqrt{2} }{ 1-\sqrt{2} }=\frac{5(1-\sqrt{2})}{1-2}=\frac{5(1-\sqrt{2})}{-1}= -5(1-\sqrt{2})\] Important: you have take the opposite sign when multiplying a denominator like that. Because then you have a remarkable product of (a+b)(a-b)=(a²-b²). So when there is an i in the denominator is like having a square root in the denominator, because i= squareroot of -1. So, you need to do what with the original expression?

OpenStudy (anonymous):

sorry the first expression should be \[\frac{ 5 }{ \sqrt{2} }\] not \[\frac{ 5 }{ ? }\]